OpenStudy (gabrielah96):

Find exact value of cos75 using half angle identities

OpenStudy (anonymous):

cos(75) = cos(150/2)

OpenStudy (anonymous):

\[\cos(150/2) = \sqrt{(1-(\sqrt{3}/2))/2}\]

OpenStudy (anonymous):

cos(75) = \[(\sqrt{2-\sqrt{3}})/2\]